# Hitting probability for Reflected Brownian Motion at Small Target

### v2

@@ 3. Main Result
In this section we define the hitting target, absorbing boundary and reflecting boundaries. We will also states our main result.
@@ 3.1. Description of Hitting Target
Let $\epsilon>0$ be a very small constant ($\epsilon\ll R$) and then our target (which is also absorbing) is $B_{\epsilon}(P)\cap\partial\Omega$, which will be denoted by $\mathrm{tg}$.
@@
Throughout the paper we assume there is a path inside $\Omega$ that connects a point in $\mathrm{tg}$ and $\mathrm{abs}$.
@@
We now give definition of the absorbing boundaries.
@@
Let $B^{x_{0}}_{t}$ be the reflected Brownian Motion in $\Omega$. Let $\tau=\inf\{t\geq 0:B^{x_{0}}_{t}\in\mathrm{abs}\cup\mathrm{tg}\}$ be the stopping time that RBM hits absorbing boundaries. We consider the stopped process $B^{x_{0}}_{\tau\wedge t}$ to be the RBM associated to the absorbing boundaries. In the later texts, without specification, $B^{x_{0}}_{t}$ denotes $B^{x_{0}}_{\tau\wedge t}$.
@@ 3.2. Main Result
Let $\Omega$ be a SULD in $\mathbb{R}^{n}$ (identify $\mathbb{C}$ as $\mathbb{R}^{2}$) and $P$ be a point on $\partial\Omega$, with target, reflecting boundary and absorbing boundary defined above. Let $(B^{x_{0}}_{t},t>0)$ be the reflected Brownian Motion starting at $x_{0}\in\Omega$, and $\tau=\inf\{t\geq 0:B^{x_{0}}_{t}\in\mathrm{abs}\}$. Define
@@
We have a stronger result, which confirm the asymptotic behavior of $\mathbb{P}(B^{z_{0}}_{\tau}\in\mathrm{tg})$ for certain domain.
@@
Let $\Omega$ be an analytic simply connected domain in $\mathbb{C}$. If in addition there is only one component of absorbing boundary, then there exists a constant $C(\Omega,z_{0})$ that depends on $\Omega$ and starting point $z_{0}$ such that
@@ 4. Standard Definitions and Standard Theorems
In this section we give some definitions to avoid confusions. Also we quote some useful standard theorems and lemmas.
Definition 4.1.
A conformal map between two domain is a holomorphic bijection between these two domains. [12] pp.206
@@
(Riemann Mapping Theorem) Let $\Omega$ be a simply connected domain and $\Omega\neq\mathbb{C}$, then there exists a conformal map between $\Omega$ and $\mathbb{D}$. [12] Chapter 8, Theorem 3.1
We also note a uniqueness result.
@@
Let $f$ be a conformal bijection between $\mathbb{D}$ and $\Omega$ where $\Omega$ is an analytic simply connected domain. Then by lemma 7.3 $f$ may be extended to a conformal bijection $g$. If in addition we fix two point on $\partial\Omega$ together with their derivatives, then such $f$ is unique.
We also have the following very useful Christoffel-Schwarz formula which maps $\mathbb{H}$ onto polygons.
Theorem 4.4.
Assume $\sum^{n-1}_{k=1}\beta_{k}<2$. Let $\mathfrak{p}$ be a polygonal region in the complex plane with vertexs $a_{1}$, $a_{2}$, …, $a_{n}$ where the angle at vertex $a_{k}$ for $k<n$ is $(1-\beta_{k})\pi$ and the angle at vertex $a_{n}$ is $(1-\beta_{n})\pi$ where
@@
Let $\Omega$ be an analytic simply connected domain, $z$ be a point inside $\Omega$ and $A$ be an Lebesgue measurable subset of $\partial\Omega$. Then by Riemann mapping theorem there exists a unique conformal bijection maps $\Omega$ onto $\mathbb{D}$ such that $f(z)=0$ and $f^{\prime}(z)>0$. Then $A$ is mapped onto $f(A)$ on $\partial\mathbb{D}$ and we define $|f(A)|$ be the Lebesgue measure of $|f(A)|$ on measurable space $\partial\mathbb{D}$. The harmonic measure of $A$ with respect to $z$, $\omega_{\Omega}(z,A)$, is then defined by $\omega_{\Omega}(z,A)=\displaystyle\frac{|f(A)|}{2\pi}$. [13]
@@
$\omega_{\Omega}(z,A)=\mathbb{P}\left(B_{\tau}^{z}\in A\right).$
[14] pp.117
@@ 5.1. Explicit Computation in $\mathbb{H}$
In this subsection we proof a result that is very similar and can be easily modified to our main result by explicitly computation in upper half plane $\mathbb{H}$.
@@
are convergent integrals, where $K$ is the complete elliptical integral of first kind.
@@
As four points $A,B,C,D$ are Lebesgue null sets we may remove them from consideration and rectangle $ABCD$ could be thought as an analytic simply connected domain. The boundaries are mapped to corresponding boundaries of the rectangle as $f_{\epsilon}$ is injective. Then by proposition 2.12 with reversed direction of $f$ (which also holds, proof is similar) we may conclude. ∎
We observe the following reflection property regarding the RBM in rectangle, which will be useful later.
@@
Let $R^{z_{0}}_{t}$ be the RBM in $ABCD$ starts at $z_{0}$ and $\tau_{B}=\inf\{t\geq 0:R^{z_{0}}_{t}\in\partial{B}\cap ABCD\}$. Let $Z^{z_{0}}_{t}=X^{x_{0}}_{t}+iY^{y_{0}}_{t}$ be the standard Brownian motion in $\mathbb{C}$ starts at $z_{0}$ such that both $X^{x_{0}}_{t}$ and $Y^{y_{0}}_{t}$ are independent standard Brownian motions on $\mathbb{R}$ start at $x_{0}$, $y_{0}$, respectively. Let $\tau^{\prime}_{B}=\inf\{t\geq 0:Z^{z_{0}}_{t}\in\partial B\}$, then with probability one, we have
@@
By taking minus sign which preserve RBM, the stopped RBM $-R^{z_{0}}_{t\wedge{\tau_{B}}}$ satisfies the stochastic differential equation in the form of (1) as it has normal reflection on $AB$. Let $H^{-z_{0}}_{t}$ be the RBM in $\mathbb{H}$ starts at $-z_{0}$ and that $\tau=\inf\{t\geq 0:H^{-z_{0}}_{t}\in\partial(-B)\cap\mathbb{H}\}$. Then stopped RBM $H^{-z_{0}}_{t\wedge\tau}$ still satisfies the stochastic differential equation in the form of (1). Thus by uniqueness stated in Theorem 2.2 we see that with probability 1, $-R^{z_{0}}_{t\wedge{\tau_{B}}}=H^{-z_{0}}_{t\wedge\tau}$. By Definition 2.9 we can write $H^{-z_{0}}_{t\wedge\tau}=X^{-x_{0}}_{t\wedge\tau^{\prime}_{-B}}+i\left|Y^{-y_{0}}_{t\wedge\tau^{\prime}_{-B}}\right|$ where both $X^{-x_{0}}_{t\wedge\tau^{\prime}_{-B}}$ and $Y^{-y_{0}}_{t\wedge\tau^{\prime}_{-B}}$ are stopped standard Brownian motion in $\mathbb{C}$. By taking minus sign back and notice that with probability 1, $-X^{-x_{0}}_{t\wedge\tau^{\prime}_{-B}}=X^{x_{0}}_{t\wedge\tau^{\prime}_{B}}$ and $-Y^{-y_{0}}_{t\wedge\tau^{\prime}_{-B}}=Y^{y_{0}}_{t\wedge\tau^{\prime}_{B}}$, we have that
@@
RBM in rectangle $ABCD$ is equivalent to the standard Brownian motion in the strip $S_{\epsilon}:=\{z\in\mathbb{C}:-L_{\epsilon}<\Re(z)<0\}$ with target at $T_{\epsilon}:=\{z\in\mathbb{C}:\Re(z)=-L_{\epsilon}\}$ and absorbing boundary at Imaginary axis.
@@
By equivalent we mean there is an bijective map $h$ between Brownian motion in strip (starts inside rectangle $ABCD$) and RBMs in the rectangle, preserving the corresponding stopping condition. The intuition is shown in figure 3: we may reflect $ABCD$ over its sides with length $L_{\epsilon}$ over and over again to get the strip we need.
@@ 5.2. Proof of Main Results in $\mathbb{C}$
Now we are ready to proof the main results by explicit computation above. We starts from stronger result.
@@
Firstly notice that ${\angle{A_{t_{1}}PA_{t_{2}}}\leq 2\pi}$ for all $t_{1},t_{2}$ which means the supremum is well-defined and denote it by $s$. We may extract a subsequence $\angle{A^{(n)}_{t_{1}}PA^{(n)}_{t_{2}}}\to s$. As $\Omega$ is an analytic simply connected domain, observe that $\partial{\Omega}\cap B_{R}(P)$ is compact so by choosing a convergence subsequence $\left(A^{(n_{k})}_{t_{1}}\right)_{k\geq 1}$ of $\left(A^{(n)}_{t_{1}}\right)_{n\geq 1}$ and another convergence subsequence $\left(A^{\left(n_{k_{j}}\right)}_{t_{2}}\right)_{j\geq 1}$ of $\left(A^{(n_{k})}_{t_{2}}\right)_{k\geq 1}$, we find $\left(A^{\left(n_{k_{j}}\right)}_{t_{1}},A^{\left(n_{k_{j}}\right)}_{t_{2}}\right)_{j\geq 1}$ converge to some $A_{t_{1}},A_{t_{2}}\in\partial{\Omega}\cap B_{R}(P)$ and $s=\angle{A_{t_{1}}PA_{t_{2}}}$. For uniqueness we just notice that if there exists distinct pairs $(A_{t_{1}},A_{t_{2}})$ and $(A^{\prime}_{t_{1}},A^{\prime}_{t_{2}})$ such that $s=\angle{A_{t_{1}}PA_{t_{2}}}=\angle{A^{\prime}_{t_{1}}PA^{\prime}_{t_{2}}}$, then by simple Euclidean Geometry at least one of angle $\angle{MPN}$ where $M,N\in\{A_{t_{1}},A_{t_{2}},A^{\prime}_{t_{1}},A^{\prime}_{t_{2}}\}$ is greater than $s$. Contradiction. ∎
@@
Let $\{C_{t}:t\in I\}$ be the collection of closure of components of $B_{P}(\epsilon)\cap\partial\Omega$ that doesn’t contain $P$. Then as connected component is closed it is compact. So $\mathrm{dist}(P,C_{t})>0$ for all $t\in I$. Now if $\inf\{\mathrm{dist}(P,C_{t}):t\in I\}=0$ then $I$ is infinite and we may take a subsequence $t_{j}$ such that $\mathrm{dist}(P,C_{t_{j}})\to 0$. Now by compactness $\mathrm{dist}(P,C_{t})$ is attended by some point $M_{t}\in C_{t}$, so for $j$ large we have all $M_{t_{j}}$ contained in a small disk $B_{P}(\delta)$ ($\delta<\epsilon$) and thus, as $C_{t_{j}}$ intersect $\partial\Omega$ at two points then their length is greater than $\epsilon-\delta$ for all $j$ large enough. But we have infinitely many such $j$ and thus $\ell(\partial\Omega)\geq\sum_{t}\ell(C_{t})=\infty$, contradiction. So $\inf\{\mathrm{dist}(P,C_{t}):t\in I\}=\xi>0$. Now Take $\epsilon<\xi$ and we are done. ∎
So it is enough to consider one component in $\mathrm{tg}$ and $\epsilon$ very small. Now we can turn to proof the main result. We firstly uniformize our domain. By Riemann Mapping Theorem there is an unique conformal bijection $h$ (extended analytically to boundary) that maps $\Omega$ onto unit disk $\mathbb{D}$ with boundary points $P$ mapped into $X$, $B$ mapped into $Y$ with non-vanishing derivative at $P$ and $B$. Let $f(C)=Z$. This is true because we assume $\Omega$ has an analytic boundary. This process is shown in figure 4.
@@
Now we can proof the stronger result. Let $f_{1}(z)$ be the extension of $f^{-1}$ beyond $\mathbb{D}$ as $\Omega$ is a analytic simply connected domain (see lemma 7.3). By open mapping theorem image of a neighborhood of $\partial\mathbb{D}$ via $f_{1}$ is an open neighborhood $\Delta$ of $\partial\Omega$. Let fixed $\delta$ so small that $B_{P}(\delta)$ is also contained in $\Delta$. Take $\epsilon<\delta$, then $f_{1}$ is invertible on $B_{P}(\epsilon)$ with holomprohic inverse $g_{1}$ that has derivatives bounded away from 00. We can get an estimation for lower bound using the univalent map $f_{1}$ which has non-vanishing derivative on $\partial\mathbb{D}$. Also, for any $\eta>0$ we may choose $\eta_{1}$ such that for any $\epsilon<\eta_{1}$, $|g_{1}^{\prime}(z)-g_{1}^{\prime}(P)|\leq\eta$ for all $z\in B_{P}(\epsilon)$. Consider when $\eta<|g^{\prime}_{1}(P)|$ and $\epsilon$ is sufficiently small, then by boundary correspondence, for all $z\in\partial g_{1}(B_{P}(\epsilon))$, $f_{1}(z)\in\partial(B_{P}(\epsilon))$. Consider the line segment joining $z$ and $g_{1}(P)$, let it be $\gamma(\xi)$. Then,
@@
There is an arc of length independent of $\epsilon$ that contained in $f(\mathrm{abs})$.
@@
Just a modification of length estimations in proof of Theorem 3.4. ∎
Observe that, if we let $c_{1}$ be an arc contained in $\mathrm{abs}$ and $c_{2}$ be another arc containing $f(\mathrm{abs})$ (for example, such $c_{2}$ can be the arc with endpoints $Y$, $Z$, which does not intersect $f(\mathrm{tg})$). Let us consider the target hitting probability for the absorbing boundary to be $c_{1}$, $f(\mathrm{abs})$ and $c_{2}$, which are $\mathbb{P}_{1}$, $\mathbb{P}_{2}$, $\mathbb{P}_{3}$, respectively, then $\mathbb{P}_{3}\leq\mathbb{P}_{2}\leq\mathbb{P}_{1}$. As by Theorem 3.4 $\mathbb{P}_{3}$ and $\mathbb{P}_{1}$ are comparable to $\frac{1}{\log\epsilon}$ as $\epsilon\downarrow 0$, $\mathbb{P}_{2}$ is as well. ∎

### v4

@@ 3. Main Result
In this section, we define the hitting target, absorbing boundary, and reflecting boundaries. We will also state our main result.
@@ 3.1. Description of Hitting Target
Let $\epsilon>0$ be a very small constant ($\epsilon=o(1)$) and then our target (which is also absorbing) is $B_{\epsilon}(P)\cap\partial\Omega$, which will be denoted by $\mathrm{tg}$.
@@
Throughout the paper, we assume there is a path inside $\Omega$ that connects a point in $\mathrm{tg}$ and $\mathrm{abs}$.
@@
We now give the definition of absorbing boundaries.
@@
Let $B^{x_{0}}_{t}$ be the reflected Brownian Motion in $\Omega$. Let $\tau=\inf\{t\geq 0:B^{x_{0}}_{t}\in\mathrm{abs}\cup\mathrm{tg}\}$ be the stopping time that RBM hits the absorbing boundary or target. We consider the stopped process $B^{x_{0}}_{\tau\wedge t}$ to be the stopped RBM associated with the boundaries. In the later texts, we use $B^{x_{0}}_{t}$ to denote $B^{x_{0}}_{\tau\wedge t}$ if there is no ambiguity.
@@ 3.2. Main Result
Let $\Omega$ be a SULD in $\mathbb{R}^{n}$ (identify $\mathbb{C}$ as $\mathbb{R}^{2}$) and $P$ be a point on $\partial\Omega$, with the target, reflecting boundary and absorbing boundary defined above in Definition 3.1. Let $(B^{x_{0}}_{t},t>0)$ be the reflected Brownian Motion starting at $x_{0}\in\Omega$, and $\tau=\inf\{t\geq 0:B^{x_{0}}_{t}\in\mathrm{abs}\}$. Define
@@
We have a stronger result, which confirms the asymptotic behavior of $\mathbb{P}(B^{z_{0}}_{\tau}\in\mathrm{tg})$ for certain domains.
@@
Let $\Omega$ be an analytic simply connected domain in $\mathbb{C}$. If, in addition, there is only one component of absorbing boundary, then there exists a constant $C(\Omega,z_{0})$ that depends on $\Omega$ and starting point $z_{0}$ such that
@@ 4. Standard Definitions and Standard Theorems
Let $f$ be a conformal bijection between $\mathbb{D}$ and $\Omega$ where $\Omega$ is an analytic simply connected domain. Then by lemma 7.3 $f$ may be extended to a conformal bijection $g$. If, in addition, we fix two points on $\partial\Omega$ together with their derivatives, then such $f$ is unique. [12]
@@
(Christoffel-Schwarz formula) Assume $\sum^{n-1}_{k=1}\beta_{k}<2$. Let $\mathfrak{p}$ be a polygonal region in the complex plane with vertexs $a_{1}$, $a_{2}$, …, $a_{n}$ where the angle at vertex $a_{k}$ for $k<n$ is $(1-\beta_{k})\pi$ and the angle at vertex $a_{n}$ is $(1-\beta_{n})\pi$ where
@@
Let $\Omega$ be an analytic simply connected domain, $z$ be a point inside $\Omega$ and $A$ be a Lebesgue measurable subset of $\partial\Omega$. Then, by the Riemann mapping theorem, there exists a unique conformal bijection maps $\Omega$ onto $\mathbb{D}$ such that $f(z)=0$ and $f^{\prime}(z)>0$. Then $A$ is mapped onto $f(A)$ on $\partial\mathbb{D}$ and we define $|f(A)|$ be the Lebesgue measure of $|f(A)|$ on measurable space $\partial\mathbb{D}$. The harmonic measure of $A$ with respect to $z$, $\omega_{\Omega}(z,A)$, is then defined by $\omega_{\Omega}(z,A)=\displaystyle\frac{|f(A)|}{2\pi}$. [12]
@@
$\omega_{\Omega}(z,A)=\mathbb{P}\left(B_{\tau}^{z}\in A\right)$
where $\tau=\inf\{t:B^{z}_{t}\in A\}$. [14] pp.117
@@ 5.1. Explicit Computation in $\mathbb{H}$
In this subsection, we prove a result that is very similar and can be easily modified to our main result by explicit computation in the upper half plane $\mathbb{H}$.
@@
are convergent integrals, where $K$ is the complete elliptical integral of the first kind.
@@
As four points $A,B,C,D$ are Lebesgue null sets, we may remove them from consideration, and rectangle $ABCD$ could be thought of as an analytic simply connected domain. The boundaries are mapped to corresponding boundaries of the rectangle as $f_{\epsilon}$ is injective. Then by proposition 2.12, we may conclude. ∎
We observe the following reflection property regarding the RBM in the rectangle, which will be useful later.
@@
Let $R^{z_{0}}_{t}$ be the RBM in $ABCD$ starts at $z_{0}$ and $\tau_{B}=\inf\{t\geq 0:R^{z_{0}}_{t}\in\partial{B}\cap ABCD\}$. Let $Z^{z_{0}}_{t}=X^{x_{0}}_{t}+iY^{y_{0}}_{t}$ be the standard Brownian motion in $\mathbb{C}$ starts at $z_{0}$ such that both $X^{x_{0}}_{t}$ and $Y^{y_{0}}_{t}$ are independent standard Brownian motions on $\mathbb{R}$ start at $x_{0}$, $y_{0}$, respectively. Let $\tau^{\prime}_{B}=\inf\{t\geq 0:Z^{z_{0}}_{t}\in\partial B\}$, then up to indistinguishability,
@@
By taking the minus sign, which preserves RBM, the stopped RBM $-R^{z_{0}}_{t\wedge{\tau_{B}}}$ satisfies the stochastic differential equation in the form of (1) as it has a normal reflection on $AB$. Let $H^{-z_{0}}_{t}$ be the RBM in $\mathbb{H}$ starts at $-z_{0}$ and that $\tau=\inf\{t\geq 0:H^{-z_{0}}_{t}\in\partial(-B)\cap\mathbb{H}\}$. Then stopped RBM $H^{-z_{0}}_{t\wedge\tau}$ still satisfies the stochastic differential equation in the form of (1). Thus by uniqueness stated in Theorem 2.2 we see that $-R^{z_{0}}_{t\wedge{\tau_{B}}}=H^{-z_{0}}_{t\wedge\tau}$ up to indistinguishability. By Definition 2.9 we can write $H^{-z_{0}}_{t\wedge\tau}=M^{-x_{0}}_{t\wedge\tau^{\prime}_{-B}}+i\left|N^{-y_{0}}_{t\wedge\tau^{\prime}_{-B}}\right|$ where both $M^{-x_{0}}_{t\wedge\tau^{\prime}_{-B}}$ and $N^{-y_{0}}_{t\wedge\tau^{\prime}_{-B}}$ are stopped standard Brownian motion in $\mathbb{C}$. By taking minus sign back and notice that with probability 1, $-M^{-x_{0}}_{t\wedge\tau^{\prime}_{-B}}:=X^{x_{0}}_{t\wedge\tau^{\prime}_{B}}$ and $-N^{-y_{0}}_{t\wedge\tau^{\prime}_{-B}}:=Y^{y_{0}}_{t\wedge\tau^{\prime}_{B}}$ are two time-changed Brownian motions in lower half plane, we have that
@@
RBM in rectangle $ABCD$ is equivalent to the standard Brownian motion in the strip $S_{\epsilon}:=\{z\in\mathbb{C}:-L_{\epsilon}<\Re(z)<0\}$ with target at $T_{\epsilon}:=\{z\in\mathbb{C}:\Re(z)=-L_{\epsilon}\}$ and absorbing boundary at Imaginary axis. By equivalent we mean that bijective map $h$ between Brownian motion in the strip (starts inside rectangle $ABCD$) and RBMs in the rectangle, preserving the corresponding stopping condition.
@@
The intuition is shown in figure 3: we may reflect $ABCD$ over its sides with length $L_{\epsilon}$ over and over again to get the strip we need.
@@ 5.2. Proof of Main Results in $\mathbb{C}$
Now, we are ready to prove the main results by explicit computation above. We start with stronger results.
@@
Firstly notice that ${\angle{A_{t_{1}}PA_{t_{2}}}\leq 2\pi}$ for all $t_{1},t_{2}$ which means the supremum is well-defined and denote it by $s$. We may extract a subsequence $\angle{A^{(n)}_{t_{1}}PA^{(n)}_{t_{2}}}\to s$. As $\Omega$ is an analytic simply connected domain, observe that $\partial{\Omega}\cap B_{R}(P)$ is compact so by choosing a convergence subsequence $\left(A^{(n_{k})}_{t_{1}}\right)_{k\geq 1}$ of $\left(A^{(n)}_{t_{1}}\right)_{n\geq 1}$ and another convergence subsequence $\left(A^{\left(n_{k_{j}}\right)}_{t_{2}}\right)_{j\geq 1}$ of $\left(A^{(n_{k})}_{t_{2}}\right)_{k\geq 1}$, we find $\left(A^{\left(n_{k_{j}}\right)}_{t_{1}},A^{\left(n_{k_{j}}\right)}_{t_{2}}\right)_{j\geq 1}$ converge to some $A_{t_{1}},A_{t_{2}}\in\partial{\Omega}\cap B_{R}(P)$ and $s=\angle{A_{t_{1}}PA_{t_{2}}}$. For uniqueness, we just notice that if there exist distinct pairs $(A_{t_{1}},A_{t_{2}})$ and $(A^{\prime}_{t_{1}},A^{\prime}_{t_{2}})$ such that $s=\angle{A_{t_{1}}PA_{t_{2}}}=\angle{A^{\prime}_{t_{1}}PA^{\prime}_{t_{2}}}$, then by simple Euclidean Geometry at least one of angle $\angle{MPN}$ where $M,N\in\{A_{t_{1}},A_{t_{2}},A^{\prime}_{t_{1}},A^{\prime}_{t_{2}}\}$ is greater than $s$. Contradiction. ∎
@@
Let $\{C_{t}:t\in I\}$ be the collection of closure of components of $B_{P}(\epsilon)\cap\partial\Omega$ that doesn’t contain $P$. Then as the connected component is closed, it is compact. So $\mathrm{dist}(P,C_{t})>0$ for all $t\in I$. Now if $\inf\{\mathrm{dist}(P,C_{t}):t\in I\}=0$ then $I$ is infinite and we may take a subsequence $t_{j}$ such that $\mathrm{dist}(P,C_{t_{j}})\to 0$. Now by compactness $\mathrm{dist}(P,C_{t})$ is attended by some point $M_{t}\in C_{t}$, so for $j$ large we have all $M_{t_{j}}$ contained in a small disk $B_{P}(\delta)$ ($\delta<\epsilon$) and thus, as $C_{t_{j}}$ intersect $\partial\Omega$ at two points then their length is greater than $\epsilon-\delta$ for all $j$ large enough. But we have infinitely many such $j$ and thus $\ell(\partial\Omega)\geq\sum_{t}\ell(C_{t})=\infty$, contradiction. So $\inf\{\mathrm{dist}(P,C_{t}):t\in I\}=\xi>0$. Now take $\epsilon<\xi$, and we are done. ∎
So it is enough to consider one component in $\mathrm{tg}$ and $\epsilon$ very small. Now, we can turn to prove of the main result. We first uniformize our domain. By Riemann Mapping Theorem, there is a unique conformal bijection $h$ (extended analytically to the boundary) that maps $\Omega$ onto unit disk $\mathbb{D}$ with boundary points $P$ mapped into $X$, $B$ mapped into $Y$ with non-vanishing derivative at $P$ and $B$. Let $f(C)=Z$. This is true because we assume $\Omega$ has an analytic boundary. This process is shown in figure 4.
@@
Now we can prove the stronger result. Let $f_{1}(z)$ be the extension of $f^{-1}$ beyond $\mathbb{D}$ as $\Omega$ is a analytic simply connected domain (see lemma 7.3). By open mapping theorem image of a neighborhood of $\partial\mathbb{D}$ via $f_{1}$ is an open neighborhood $\Delta$ of $\partial\Omega$. Let fixed $\delta$ so small that $B_{P}(\delta)$ is also contained in $\Delta$. Take $\epsilon<\delta$, then $f_{1}$ is invertible on $B_{P}(\epsilon)$ with holomprohic inverse $g_{1}$ that has derivatives bounded away from 00. We can get an estimation for the lower bound using the univalent map $f_{1}$, which has a non-vanishing derivative on $\partial\mathbb{D}$. Also, for any $\eta>0$ we may choose $\eta_{1}$ such that for any $\epsilon<\eta_{1}$, $|g_{1}^{\prime}(z)-g_{1}^{\prime}(P)|\leq\eta$ for all $z\in B_{P}(\epsilon)$. Consider when $\eta<|g^{\prime}_{1}(P)|$ and $\epsilon$ is sufficiently small, then by boundary correspondence, for all $z\in\partial g_{1}(B_{P}(\epsilon))$, $f_{1}(z)\in\partial(B_{P}(\epsilon))$. Consider the line segment joining $z$ and $g_{1}(P)$, let it be $\gamma(\xi)$. Then,
@@
There is an arc of length independent of $\epsilon$ that is contained in $f(\mathrm{abs})$.
@@
Just a modification of length estimations in proof of Lemma 5.7. ∎
Observe that if we let $c_{1}$ be an arc contained in $\mathrm{abs}$ and $c_{2}$ be another arc containing $f(\mathrm{abs})$ (for example, such $c_{2}$ can be the arc with endpoints $Y$, $Z$, which does not intersect $f(\mathrm{tg})$). Let us consider the target hitting probability for the absorbing boundary to be $c_{1}$, $f(\mathrm{abs})$ and $c_{2}$, which are $\mathbb{P}_{1}$, $\mathbb{P}_{2}$, $\mathbb{P}_{3}$, respectively, then $\mathbb{P}_{3}\leq\mathbb{P}_{2}\leq\mathbb{P}_{1}$. As by Theorem 3.4 $\mathbb{P}_{3}$ and $\mathbb{P}_{1}$ are comparable to $\frac{1}{\log\epsilon}$ as $\epsilon\downarrow 0$, $\mathbb{P}_{2}$ is as well. ∎